### v1

# Optimal Rank-Metric Codes with Rank-Locality from Drinfeld Modules

Abstract.
We introduce a new technique to construct rank-metric codes using the arithmetic theory of Drinfeld modules over global fields, and Dirichlet Theorem on polynomial arithmetic progressions. Using our methods, we obtain a new infinite family of optimal rank-metric codes with rank-locality, i.e. every code in our family achieves the information theoretical bound for rank-metric codes with rank-locality.
2020 Mathematics Subject Classification:
11T06,11T71,94B27
1. Introduction
Let $q$ be a prime power and $m$ a positive integer. Rank-metric codes are an important family of error correcting codes that are used for reliable communication over noisy channels [1, 5, 12, 18, 23] and for cryptography [7, 10, 21]. They have been widely studied in the literature due to their interesting algebraic combinatorial properties that allow to encode information in a way that the obtained codeword are resilient to rank errors, i.e. the noise is modeled as a sum of a low rank matrix.
Locality is a foundational concept in information theory [22, 25, 26]. It allows for efficient recovery of encoded information that is partially lost in distributed storage systems. More specifically, if the information vector is encoded and distributed over multiple servers, locality allows to recover the information contained in a single server (that might be unavailable for various reasons) using only a few other servers.
Rank-metric codes with rank-locality are an important subclass of rank-metric codes that allow efficient corrections of criss-cross failures in data centers (for additional detail on this see for example [14, Example 1]). Various constructions with flexible parameters are available for locally recoverable codes over the Hamming metric [4, 3, 6, 8, 11, 15, 16, 17] but only one for the rank-metric codes [14] (with very restrictive parameters). The difficulty that arises in the rank-metric context is that classical constructions fail due to the non-commutative nature of the message space, which naturally consists of $q$-linearized polynomials with coefficients in ${\mathbb{F}}_{q^{m}}$.
This paper focuses on general constructions of optimal rank-metric codes with rank-locality, i.e. codes whose dimension achieves the information theoretical limits imposed by the requiring locality and minimum distance. Our methods leverage powerful tools from the theory of Drinfeld modules. The shape of our message space is drawn by the decomposition of Anderson motives as free modules over ${\mathbb{F}}_{q^{m}}[T]$ (using the natural Drinfeld action). The evaluation points that provide the encoding are picked by carefully selecting an $h\in{\mathbb{F}}_{q}[T]$ and a Drinfeld module $\phi$ so that the $h$-torsion is defined of ${\mathbb{F}}_{q^{m}}$. Differently from [22] (where the authors get classical locality, as explained in [14, Section 3]), we actually get proper rank-locality, as required in [14]. Guaranteeing the existence of such $h$, relies on an effective version of Dirichlet Theorem on polynomial arithmetic progression, which we derive from the $L$-series technique in [20, Chapter 4].
For the rest of this introduction we describe our main information theoretical result. The reader who is not familiar with the rank-metric code material should refer to section 2.3 before diving into the coding theory result below. Our main result is the following.
Theorem 1.1.
Let $q$ be a prime power, and let $m\geq 4$ be an integer. Let $R,\ell$ be positive integers such that
$\ell R<\frac{m}{2}-\frac{\log(m)}{\log(q)}-\frac{\log\left(5\right)}{\log(q)},$
$2\leq\delta<R$, $s+1\leq\ell<q$, and $q$ is coprime to $R$. Then, there exists a rank-metric code in ${\mathbb{F}}_{q}^{m\times\ell R}$, rank-locality $(r,\delta)$, $R=r+\delta-1$, and ${\mathbb{F}}_{q}$-dimension equal to $m(s+1)r$. Moreover, this code has rank distance $\ell R-Rs-r+1$, thus it is optimal.
Notice that most of these condition are non restrictive, but simply guarantee that there is enough space in ${\mathbb{F}}_{q}^{m\times\ell R}$ to build a optimal code with rank-locality having non-trivial parameters. For example, the code lives in a space of dimension $m\ell R$ and its ${\mathbb{F}}_{q}$-dimension will be $m(s+1)r$, so it is expected to have the inequality $s+1\leq\ell$, or for example $\delta\geq 2$ is necessary to have rank-locality and $\delta\leq R-1$ simply requires that the distance of the local code (i.e. the code in ${\mathbb{F}}_{q}^{m\times R}$ where one punctures at all columns except the ones in the locality set) makes sense.
2. Preliminaries
2.1. Drinfeld Modules
Throughout the article, let $q$ be a prime power and ${\mathbb{F}}_{q}$ the finite field with $q$ elements. We denote by ${\mathbb{F}}_{q}[T]$ the ring of polynomials over ${\mathbb{F}}_{q}$ and by ${\mathbb{F}}_{q}(T)$ the rational function field in the variable $T$. A function field $F$ over ${\mathbb{F}}_{q}$ is denoted by $F/{\mathbb{F}}_{q}$, or simply $F$ whenever the constant field is understood. Let $P$ be a place of $F$. We denote by $\mathcal{O}_{P}$ the valuation ring at $P$. We will also use other standard notation for function fields, for which we refer to [24].
Definition 2.1.
Let ${\mathbb{F}}_{q^{n}}$ be a finite extension of ${\mathbb{F}}_{q}$. We define the norm
$\operatorname{Nr}:{\mathbb{F}}_{q^{n}}$ $\longrightarrow{\mathbb{F}}_{q^{n}}$
$\alpha$ $\longmapsto\prod_{i=0}^{n-1}\alpha^{q^{i}}$
Definition 2.2.
Let $P\in{\mathbb{F}}_{q}[T]$ be an irreducible polynomial and $\mathfrak{p}=(P)$ be the ideal generated by $P$. We say ${\mathbb{F}}_{q}[T]/\mathfrak{p}$ is the residue field of $\mathfrak{p}$ and we shall denote it by ${\mathbb{F}}_{\mathfrak{p}}$.
In this section, we provide some background about the theory of Drinfeld modules. For a comprehensive treatise on the subject, refer to [19]. Throughout this section we assume $K\supseteq{\mathbb{F}}_{q}$ is a field and $a\in{\mathbb{F}}_{q}[T]$.
Definition 2.3.
Let $n\geq 0$. A polynomial $f(x)=\sum_{i=0}^{n}a_{i}x^{q^{i}}$ in $K[x]$ is called a $q$-linearized polynomial. If $\operatorname{deg}(f)=q^{n}$, we say that $f$ has $q$-degree equal to $n$ and we denote $q^{i}$ by $[i]$. The ring of $q$-linearized polynomials is denoted by $K\langle x\rangle$, and in general, it is a noncommutative ring under addition and composition.
Definition 2.4.
Let $K$ be a commutative ${\mathbb{F}}_{q}$-algebra and $\tau$ be an indeterminate. Let $K\{\tau\}$ be the set of polynomials of the form $\sum_{i=0}^{n}a_{i}\tau^{i}$ such that $a_{i}\in K$ and $n\geq 0$. In $K\{\tau\}$, for $\{a_{i},b_{i}\}_{i=0,\ldots,n}\in K$, we define the addition as the usual addition of polynomials
$\sum_{i=0}^{n}a_{i}\tau^{i}+\sum_{i=0}^{n}b_{i}\tau^{i}=\sum_{i=0}^{n}(a_{i}+b_{i})\tau^{i}$
while the multiplication is defined as follows for any $a,b\in K$
(2.1) $(a\tau^{i})(b\tau^{j}):=ab^{q^{i}}\tau^{i+j}.$
With these operations $K\{\tau\}$ is a noncommutative ring called the ring of twisted polynomials.
Definition 2.5.
Let $K$ be an ${\mathbb{F}}_{q}[T]$-field, (i.e. a field equipped with an ${\mathbb{F}}_{q}$-algebra homomorphism $\gamma:{\mathbb{F}}_{q}[T]\rightarrow K$). A Drinfeld module of rank $R\geq 1$ is an ${\mathbb{F}}_{q}$-algebra homomorphism
$\phi:{\mathbb{F}}_{q}[T]\longrightarrow$ $K\{\tau\}$
$a\longmapsto$ $\phi_{a}=\gamma(a)+\sum^{n}_{i=1}h_{i}(a)\tau^{i}$
where $h_{n}(a)\neq 0$ and $n=\operatorname{deg}_{T}(a)\cdot R$.
The morphism $\gamma$ is called the structure morphism of $\phi$. Observe that for any $a=\sum_{i=0}^{n}a_{i}T^{i}$, we have $\phi_{a}=\sum_{i=0}^{n}a_{i}\phi_{T}^{i}$. In other words, $\phi_{T}$ determines $\phi$ entirely. The Carlitz module is the Drinfeld module of rank $1$ and defined by $\phi_{T}=\gamma(T)+\tau$ [19, Definition 3.2.5]. Moreover, $K\{\tau\}$ is related to the ring of $q$-linearized polynomials $K\langle x\rangle$ via the following isomorphism
$\iota:K\{\tau\}$ $\longrightarrow K\langle x\rangle$
$a_{i}\tau^{i}$ $\longmapsto a_{i}x^{q^{i}}.$
Definition 2.6.
We say that $\iota(\phi_{a})$ is the $a$-division polynomial of $\phi$ and denote it by $\phi_{a}(x)$.
It is a well-known fact in finite field theory that a $q$-linearized polynomial in $K\langle x\rangle$ is also an ${\mathbb{F}}_{q}$-linear operator on $K$ (see [19, Chapter 3.1]). Therefore, one can talk about $\operatorname{ker}\left(\phi_{a}(x)\right)\subseteq\overline{K}$.
Definition 2.7.
An element $\alpha\in\operatorname{Ker}\left(\phi_{a}(x)\right)\subseteq\overline{K}$ is called an $a$-torsion point. The set of all $a$-torsion points is denoted by $\phi[a]$. In addition, we denote by $K(\phi[a])$ the splitting field of $\phi_{a}(x)$ over $K$ and refer to it as the $a$-division field of $\phi$.
It turns out that the set of $a$-torsion points has a more general structure than just an ${\mathbb{F}}_{q}$-vector space.
Proposition 2.8.
For any fixed $a\in{\mathbb{F}}_{q}[T]$, the set of $a$-torsion points $\phi[a]$ is a finite ${\mathbb{F}}_{q}[T]-$module with scalar multiplication defined by $b\circ\alpha=\phi_{b}(\alpha)$, for some $b\in{\mathbb{F}}_{q}[T]$ and $\alpha\in\phi[a]$.
Another useful fact about the structure of the $ab$-torsion points for coprime $a,b\in{\mathbb{F}}_{q}[T]$ is the following proposition.
Proposition 2.9.
[19, Lemma 3.5.1] If $a,b\in{\mathbb{F}}_{q}[T]$ are relatively prime nonzero elements, then, $\phi[ab]=\phi[a]\times\phi[b]$ as ${\mathbb{F}}_{q}[T]$-modules.
For $K={\mathbb{F}}_{q}(T)$, we define good reduction at a prime ideal $\mathfrak{p}=(P)$ of ${\mathbb{F}}_{q}[T]$. But first, we need to recall the definition of non-Archimedean valuation on ${\mathbb{F}}_{q}(T)$.
Definition 2.10.
Let $P$ be a monic irreducible polynomial in ${\mathbb{F}}_{q}[T]$ and $\mathfrak{p}$ the ideal generated by P. For any nonzero $h\in{\mathbb{F}}_{q}(T)$, we can write $h=P^{n}\cdot\frac{f}{g}$ for some $n\in\mathbb{Z}$ with $f,g\in{\mathbb{F}}_{q}[T]$, $\operatorname{gcd}(f,g)=1$ and $g\neq 0$ such that $P\nmid fg$. The discrete valuation $v_{\mathfrak{p}}:{\mathbb{F}}_{q}(T)\rightarrow\mathbb{Q}\cup\{\infty\}$ is defined as
$v_{\mathfrak{p}}(h)=\begin{cases}n&\text{ if }h\neq 0\\ \infty&\text{ if }h=0\\ \end{cases}$
Definition 2.11.
[19, Remark 7.1.18] Let $\mathfrak{p}$ be a prime ideal in ${\mathbb{F}}_{q}[T]$ and ${\mathbb{F}}_{\mathfrak{p}}$ be the residue field at $\mathfrak{p}$. Let $\phi:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{q}(T)\left\{\tau\right\}$ be a Drinfeld module of rank $R$. Let $\phi_{T}=\sum_{i=0}^{R}a_{i}\tau^{i}$. We say that $\phi$ has a good reduction at $\mathfrak{p}$ if it satisfies the following conditions:
  • $v_{\mathfrak{p}}(a_{R})=0$,
  • $v_{\mathfrak{p}}(a_{i})\geq 0$ for $10\leq i\leq R-$ 1.
If these conditions are verified, then $\overline{\phi}:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{\mathfrak{p}}\{\tau\}$ is given by reducing the coefficients of $\phi_{T}$ at $\mathfrak{p}$ and naturally extending the homomorphism using the algebra structure of ${\mathbb{F}}_{q}[T]$. In other words, $\overline{\phi}$ is a Drinfeld module of rank $R$ over ${\mathbb{F}}_{\mathfrak{p}}$.
It is immediate to see that if $\phi$ in the preceding definition is a Carlitz module, then it has good reduction everywhere. Next, we present the following theorem about a family of Drinfeld modules of Rank $1$.
Theorem 2.12.
Let $P,h$ be polynomials in ${\mathbb{F}}_{q}[T]$ such that $P$ is monic irreducible with a residue field ${\mathbb{F}}_{\mathfrak{p}}$, and $g$ be a nonzero element in ${\mathbb{F}}_{q}(T)$ such that $v_{\mathfrak{p}}(g)=v_{\mathfrak{p}}(h)=0$. Denote by $\overline{g}\in{\mathbb{F}}_{\mathfrak{p}}$ the reduction of $g$ modulo $P$. Let $\phi^{(g)}:{\mathbb{F}}_{q}[T]\rightarrow{\mathbb{F}}_{q}(T)\{\tau\}$ be a Drinfeld module of rank $1$ where $\phi_{T}^{(g)}=T+g\tau$. Then, $P$ splits completely in the $h$-division field of $\phi$ if and only if $P\equiv\operatorname{Nr}_{{\mathbb{F}}_{\mathfrak{p}}/{\mathbb{F}}_{q}}(\overline{g})\mod h$.
Proof.
See [19, Theorem 7.1.19]
Corollary 2.13.
Let $R,\ell\geq 1$. Let $\{a_{1},\ldots,a_{\ell}\}\subseteq{\mathbb{F}}_{q}^{*}$. Let $h=\prod_{i=1}^{\ell}(T^{R}-a_{i})$ and $P\equiv 1\mod h$ be a monic irreducible polynomial in ${\mathbb{F}}_{q}[T]$ with residue field ${\mathbb{F}}_{q^{m}}$. Let $\phi^{(1)}$ be the Drinfeld module of rank $1$ defined in Theorem 2.12 with $g=1$ and $\phi_{T}^{(1)}=T+\tau$. Then, the $h$-division polynomial of $\bar{\phi}^{(1)}$ splits completely in ${\mathbb{F}}_{q^{m}}$.
Proof.
Observe that $g=1$, and from the choice of $P$ we get immediately that $P$ is a totally split place for the $h$-division field $M={\mathbb{F}}_{q}(T)(\phi^{(1)}[h])$. This means that all places of ${\mathbb{F}}_{q}(T)(\phi^{(1)}[h])$ lying above $P$ have relative degree $1$ over the residue field at $P$. In other words, if $\alpha\in{\mathbb{F}}_{q^{m}}$ is a root of $P$, and $R$ is any place above $P$, then $[\mathcal{O}_{R}/R:{\mathbb{F}}_{q}(\alpha)]=1$. Now we want to see that this leads to a complete split of $\bar{\phi}^{(1)}_{h}(x)$ in ${\mathbb{F}}_{q}(\alpha)[x]\cong{\mathbb{F}}_{q^{m}}[x]$. Now let $\beta$ be a root of $\phi^{(1)}_{h}(x)$ in $M$ and consider the intermediate extension $L={\mathbb{F}}_{q}(T)(\beta)$. Then we have $M\supseteq L\supseteq{\mathbb{F}}_{q}(T)$. If $P$ splits completely in $M:{\mathbb{F}}_{q}(T)$, then $P$ splits completely in $L:{\mathbb{F}}_{q}(T)$ since ramification and relative degrees are multiplicative. So we have that $\bar{\phi}^{(1)}_{h}(x)\mod P$ cannot have irreducible factors of degree larger than $1$ because of Kummer Theorem [24, Theorem 3.3.7]. It remains to show that no root of $\bar{\phi}^{(1)}_{h}(x)\mod P$ has multiplicity larger than $1$, so that the number of irreducible factors of $\bar{\phi}^{(1)}_{h}\mod P$ is exactly $\deg(\phi^{(1)}_{h}(x))=q^{R\ell}$, which forces that $\phi^{(1)}_{h}(x)$ to be totally split, as wanted. For $\beta\in\overline{\mathbb{F}}_{q}$ let us denote by $m_{\beta}$ the minimal polynomial of $\beta$ over ${\mathbb{F}}_{q}$. To see that $\bar{\phi}^{(1)}_{h}\mod P$ has no repeated roots, observe that the only specializations $\beta\in\overline{\mathbb{F}}_{q}$ of $T$ which lead to a repeated root of $\bar{\phi}^{(1)}_{h}(x)\mod m_{\beta}$ are in correspondence with the zeroes in $\overline{{\mathbb{F}}}_{q}$ of the coefficient of $x$ in $\phi^{(1)}_{h}(x)$ (simply because $\phi^{(1)}_{h}(x)$ is $q$-linearized). But, by direct computation, one sees that this coefficient is $h(T)$, and $P$ is coprime with $h$. This forces that $\alpha$ is never a specialization that leads to a repeated root of $\phi^{(1)}_{h}(x)$ or, more precisely, $\bar{\phi}^{(1)}_{h}\mod P$ is squarefree, and therefore totally split (since each of its irreducible factors has degree $1$). ∎
Observe that $K\{\tau\}$ can be seen as module over $K[T]$. For any $n\in{\mathbb{N}}$, we denote by $K\{\tau\}_{\leq n}$ the set $\{\sum^{n}_{i=0}a_{i}\tau^{i}\;|\;a_{i}\in K\}$.
Proposition 2.14.
$K\{\tau\}$ is free module of rank $R$ over $K[T]$.
Proof.
It follows from [19, Lemma 3.4.4]
2.2. $L$-series
Let $\mathbb{C}$ be the field of complex numbers, $q$ be a prime power, $A={\mathbb{F}}_{q}[T]$ be the ring of polynomials over the finite field with $q$ elements, and $h$ be a fixed element of $A$ of positive degree. Unless otherwise specified, we write $P$ to denote a monic irreducible polynomial in $A$, so a sum whose index is $P$ is meant to run only over irreducible polynomials.
Theorem 2.15.
Let $S_{m}$ denote the set of monic irreducible polynomials in $A$ of degree $m$. Then,
|#$\left|\#S_{m}-\frac{q^{m}}{m}\right|\leq 2\frac{q^{\frac{m}{2}}}{m}.$
Proof.
It follows directly from [20, Theorem 2.2]. ∎
Definition 2.16.
A Dirichlet character modulo $h$ is a function $\chi:A\rightarrow\mathbb{C}$ such that:
  • (a)
    $\chi(a+bh)=\chi(a)$ for all $a,b\in A$
  • (b)
    $\chi(a)\chi(b)=\chi(ab)$ for all $a,b\in A$
  • (c)
    $\chi(a)\neq 0$ if and only if $\gcd(a,h)=1$
From part (b), it follows immediately that $\chi(1)=1$. Generally, it can be shown that the value of a character is either zero or a root of unity, and so $|\chi(a)|\in\{0,1\}$ for every $a\in A$.
Let $\Phi(h)$ be the number of non-zero polynomials of degree less than $\deg(h)$ and relatively prime to $h$. It can be shown that there are exactly $\Phi(h)$ Dirichlet characters modulo $h$ (see [20, Chapter 4]).
Definition 2.17.
For every $a\in A$, the trivial Dirichlet character modulo $h$ is denoted by $\chi_{o}$ and defined by $\chi_{o}(a)=1$ if $\gcd(a,h)=1$, $\chi_{o}(a)=0$ if $\gcd(a,h)\neq 1$.
Let $X_{h}$ denote the set of all the Dirichlet characters modulo $h$. For every $\chi,\psi\in X_{h}$, we can define the product by $\chi\psi(a)\coloneqq\chi(a)\psi(a)$ and the inverse by $\chi^{-1}(a)\coloneqq\chi(a)^{-1}$. With these operations, $X_{h}$ can be seen as a group with identity element $\chi_{o}$. Finally, since characters are complex-value functions, we write $\overline{\chi(a)}$ or $\bar{\chi}(a)$ for the complex conjugate of $\chi(a)$. Next, we outline a result about orthogonality relations between characters.
Proposition 2.18.
[20, Proposition 4.2] Let $\chi,\psi$ be Dirichlet characters modulo $h$ and $a,b\in A$ that are coprime to $h$. Let $\delta$ be the Kronecker delta. Then,
  • (1)
    $\sum_{a}\chi(a)\overline{\psi(a)}=\Phi(h)\delta(\phi,\psi)$
  • (2)
    $\sum_{\chi}\chi(a)\overline{\chi(b)}=\Phi(h)\delta(a,b)$
[…]
From now on, for every $f\in A$, let $|f|=q^{\deg(f)}$ and $s$ be a complex variable with real part greater than 1.
Definition 2.19.
Let $\chi$ be a Dirichlet character modulo $h$. The Dirichlet L-series corresponding to $\chi$ is defined by
$L(s,\chi)=\sum_{f\text{ monic}}\frac{\chi(f)}{|f|^{s}}.$
Proposition 2.20 (Euler product).
The following equality holds
(2.2) $L(s,\chi)=\prod_{P}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}.$
In particular, for the $L$-series corresponding to the trivial character we have
(2.3) $L(s,\chi_{o})=\prod_{P|h}\left(1-\frac{1}{|P|^{s}}\right)\zeta_{A}(s)$
where
(2.4) $\zeta_{A}(s)=\sum_{\begin{subarray}{c}f\in A\\ f\text{ monic}\end{subarray}}\frac{1}{|f|^{s}}=\frac{1}{1-q^{1-s}}$
is the Riemann Zeta function for $A$.
Proof.
See [20, Chapter 4]. The proof follows from the definition of L-series and comparison with the properties of $\zeta_{A}(s)$, written in the product form
$\zeta_{A}(s)=\prod_{P}\left(1-\frac{1}{|P|^{s}}\right)^{-1}.$
Notice that, from part (c) of Definition 2.16, we can rearrange Equation (2.2) as follows:
(2.5) $\begin{split}L(s,\chi)&=\prod_{P}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}=\prod_{P\nmid h}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}\prod_{P|h}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}\\ &=\prod_{P\nmid h}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}.\end{split}$
Proposition 2.21.
[20, Propostion 4.3] Let $\chi$ be a non-trivial Dirichlet character modulo $h$. Then, $L(s,\chi)$ is a polynomial in $q^{-s}$ of degree at most $\deg(h)-1$.
2.3. Rank Metric Codes
In this section, we provide some background about rank-metric codes based on [13, Ch II.11]. Throughout this section, we denote by ${\mathbb{F}}^{m\times n}_{q}$ the $m\times n$ matrices with entries in ${\mathbb{F}}_{q}$. Now, we define the rank distance on ${\mathbb{F}}^{m\times n}_{q}$, which is of fundamental importance in the theory of rank-metric codes.
Definition 2.22.
Let $A,B\in{\mathbb{F}}^{m\times n}_{q}$ and let $\operatorname{rk}_{q}(A)$ be the usual rank of $A$. We define the rank distance between $A$ and $B$ as $d_{R}(A,B)=\operatorname{rk}_{q}(A-B)$.
One can check that this is indeed a distance on ${\mathbb{F}}^{m\times n}_{q}$.
Definition 2.23.
A (matrix) rank-metric code $\mathcal{C}$ is an ${\mathbb{F}}_{q}-$ linear subspace of ${\mathbb{F}}^{m\times n}_{q}$. We define the minimum distance or simply the distance of a nonzero code to be
$d_{R}(\mathcal{C}):=\min\{d_{R}(A,B):A,B\in\mathcal{C}\textup{ and }A\neq B\}$
For any $x=(x_{1},\ldots,x_{n})\in{\mathbb{F}}^{n}_{q^{m}}$, we can define a matrix $M_{x}\in{\mathbb{F}}^{m\times n}_{q}$ associated with $x$ up to the choice of the basis as follows. Fix a basis =$\mathcal{B}=\{b_{1},\ldots,b_{m}\}$ of ${\mathbb{F}}_{q^{m}}$ over ${\mathbb{F}}_{q}$. Then, $x_{j}=\sum_{i=1}^{m}c_{ij}b_{i}$ for some $c_{ij}\in{\mathbb{F}}_{q}$. Thus, $M_{x}=[c_{ij}]_{\begin{subarray}{c}1\leq i\leq m\\ 1\leq j\leq n\end{subarray}}$. In conclusion, we can define another type of rank-metric codes; namely (vector) rank-metric codes.
Definition 2.24.
A (vector) rank-metric code $\mathcal{C}$ is an ${\mathbb{F}}_{q^{m}}$-subspace of ${\mathbb{F}}^{n}_{q^{m}}$. The minimum distance of a nonzero code $\mathcal{C}$ is $d_{R}(\mathcal{C}):=\min\{\operatorname{rk}_{q}(M_{x}):x\in\mathcal{C}\textup{and }x\neq 0\}$.
It is immediate to see that every (vector) rank-metric code in ${\mathbb{F}}^{n}_{q^{m}}$ can be seen as a (matrix) rank-metric code in ${\mathbb{F}}_{q}^{m\times n}$ by fixing an ${\mathbb{F}}_{q}$-basis for ${\mathbb{F}}_{q^{m}}$. Note that our main construction is a (vector) rank-metric and hence also a (matrix) rank-metric code. Next, we introduce the Singleton bound for rank-metric codes.
Proposition 2.25.
[13, Theorem 11.3.5, Remark 11.3.6] Let $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{m\times n}$. Every rank-metric code with minimum distance $d_{R}(\mathcal{C})$ satisfies the following upper bound.
$|\mathcal{C}|\leq q^{\max\{m,n\}\left(\min\{m,n\}-d_{R}(\mathcal{C})+1\right)}.$
In particular, if $\mathcal{C}\subseteq{\mathbb{F}}_{q^{m}}^{n}$ is a vector rank-metric code, then,
$\dim_{{\mathbb{F}}_{q^{m}}}(\mathcal{C})\leq n-d_{R}(\mathcal{C})+1.$
Definition 2.26.
If a rank-metric code achieves equality in the previous upper bound, we call it a maximum rank-distance code (MRD).
From now on, all our rank-metric codes will be ${\mathbb{F}}_{q}$-subspaces. An important family of MRD codes is the family of Gabidulin codes (see [9]). We will need properties of these codes in what follows.
Definition 2.27.
[…]
$\operatorname{enc}:{\mathbb{F}}^{k}_{q^{m}}\longrightarrow$ ${\mathbb{F}}^{n}_{q^{m}}$
$\bm{m}\longmapsto$$\{G_{\bm{m}}(\gamma),\gamma\in\mathcal{P}\}.$
An $(n,k)$-Gabidulin Code $\mathcal{C}_{\operatorname{Gab}}$ over ${\mathbb{F}}_{q^{m}}$ is defined as
$\mathcal{C}_{\operatorname{Gab}}=\{G_{\bm{m}}(\gamma):\gamma\in\mathcal{P},\bm{m}\in{\mathbb{F}}^{k}_{q^{m}}\}.$
Definition 2.28.
[14, Definition 3] We say that $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{m\times n}$ has rank-locality $(r,\delta)$ if, for every column index $i\in\{1,\dots,n\}$, there exists a set of columns $\Gamma(i)\subseteq\{1,\dots,n\}$ such that
  • $i\in\Gamma(i)$
  • $|\Gamma(i)|\leq r+\delta-1$
  • $d_{R}(\mathcal{C}\mid_{\Gamma(i)})\geq\delta$
where $\mathcal{C}\mid_{\Gamma(i)}$ is the code obtained from all the (matrix) codewords in $\mathcal{C}$ restricted to the columns indexed by $\Gamma(i)$. The code $\mathcal{C}\mid_{\Gamma(i)}$ is said to be the local code associated with the $i$-th column. A rank-metric code $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{m\times n}$ of ${\mathbb{F}}_{q}$-dimension $K$ is said to be an $(m\times n,K)$-code. In particular, an $(n,k)$-Gabidulin code over ${\mathbb{F}}_{q^{m}}$ is an $(m\times n,mk)$ rank-metric code over ${\mathbb{F}}_{q}$. An $(m\times n,K)$-code with minimum rank distance $d_{R}$ and $(r,\delta)$ rank-locality is denoted as an $(m\times n,K,d_{R},r,\delta)$ rank-metric code.
Proposition 2.29.
[14, Equation 9] The rank distance $d_{R}(\mathcal{C})$ of an $(m\times n,mk)$ rank-metric code with $(r,\delta)$ rank-locality satisfies the following
$d_{R}(\mathcal{C})\leq n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$
Proof.
See [14, Theorem 1]
3. Optimal Rank Matric Codes with Locality from Drinfeld Modules
3.1. The construction
For this section and the theorems within, we will use the following notation so we will be able to make the statements of each theorem lighter. Let $K={\mathbb{F}}_{q}(T)$, and $\phi$ be the Carlitz module defined by $\phi_{T}=T+\tau$. Let $r\geq 1$, $\delta\geq 2$ be positive integers such that $R=r+\delta-1>0$. Let $a_{1},\dots a_{\ell}\in{\mathbb{F}}_{q}^{*}$ and $h=\prod^{\ell}_{i=1}(T^{R}-a_{i})$ for some $\ell>0$. Note that if $\gcd(q,R)\neq 1$ then $h$ is not separable. Let $P$ be a monic irreducible polynomial of degree $m$ such that $P\equiv 1\pmod{h}$. Usually, it is possible to find such polynomial even for small $m$ (see for example Section 4.1). Nevertheless, in Section 5 we will use Dirichlet Theorem for polynomial arithmetic progressions to provide a proof that there exists a wide range of parameters where we can guarantee the existence of an irreducible $P$ within the wanted residue class.
Proposition 3.1.
Let $\alpha\in{\mathbb{F}}_{q^{m}}$ be a root of $P\in{\mathbb{F}}_{q}[T]$. In this case, $\phi$ has a good reduction at $P$. Suppose $\overline{\phi}$ is the reduction of $\phi$ modulo $P$, then, the $h$-division field of $\overline{\phi}$ is ${\mathbb{F}}_{q^{m}}$. In particular, for every $i\in\{1,\dots,\ell\}$, the $(T^{R}-a_{i})$-division field of $\overline{\phi}$ is ${\mathbb{F}}_{q^{m}}$.
Proof.
By Definition 2.11, $\phi$ has good reduction modulo $P$. Consequently, $\overline{\phi}$ is a Carlitz module with $\overline{\phi}_{T}=\alpha+\tau$. Finally, by Corollary 2.13, the $h$-division polynomial of $\overline{\phi}$ splits in ${\mathbb{F}}_{q^{m}}$ which implies that for any $i\in\{1,\ldots,\ell\}$, the $(T^{R}-a_{i})$- division polynomial of $\overline{\phi}$ splits in ${\mathbb{F}}_{q^{m}}$. ∎
We are now ready to construct the code. Assume that the following conditions hold:
  1. (C1)
    $\operatorname{gcd}(q,R)=1$,
  2. (C2)
    $(s+1)r<\ell R$,
  3. (C3)
    $s+1\leq\ell<q$,
  4. (C4)
    $\ell R\leq m$.
Roughly, these conditions simply allow for enough space in ${\mathbb{F}}_{q}^{m\times\ell R}$ to construct a code with the wanted dimension and $(r,\delta)$-rank-locality.
The message space $\mathcal{M}$ is given by
=$\mathcal{M}=\left\{\sum^{s}_{k=0}g_{k}(\tau){\overline{\phi}}_{T^{R}}^{k}\mid g_{k}\in{\mathbb{F}}_{q^{m}}\{\tau\}_{\leq r-1}\right\}.$
Let $W_{i}=\overline{\phi}[T^{R}-a_{i}]$ and let $\{\beta^{(i)}_{1},\dots,\beta^{(i)}_{R}\}$ be a basis for $W_{i}$. The encoding map is given by
$\operatorname{enc}:\mathcal{M}$ $\longrightarrow{\mathbb{F}}^{\ell R}_{q^{m}}$
$f$ $\longmapsto\left(f(\beta^{(i)}_{j})\right)_{\begin{subarray}{c}i\in\{1,\dots,\ell\}\\ j\in\{1,\dots,R\}\end{subarray}}.$
Note that the encoding is injective if condition C2 is satisfied (simply because of the degree of $f$), and since the $\beta_{j}^{(i)}$’s are element of ${\mathbb{F}}_{q^{m}}$, we also need condition C4 to be able to evaluate at independent elements.
Lemma 3.2.
For any $i\in\{1,\ldots,\ell\}$ we have $\dim_{{\mathbb{F}}_{q}}\overline{\phi}[T^{R}-a_{i}]=R$.
Proof.
Note that $\overline{\phi}_{T^{R}}(x)-a_{i}x=x^{[R]}+\ldots+(\alpha^{R}-a_{i})x$. If $\alpha^{R}=a_{i}$, then, $P$ divides $T^{R}-a_{i}$ which is impossible as $\gcd(T^{R}-a_{i},P)=1$. Hence, $\overline{\phi}_{T^{R}}(x)-a_{i}x$ is separable and the result follows. ∎
Lemma 3.3.
Let
$f=\sum_{k=0}^{s}g_{k}(\tau)\overline{\phi}^{k}_{T^{R}}$
be a nonzero twisted polynomial in ${\mathbb{F}}_{q^{m}}\{\tau\}$ with $g_{k}(\tau)\in{\mathbb{F}}_{q^{m}}\{\tau\}_{\leq r-1}$. If condition C3 is satisfied (i.e. $s+1\leq\ell$), then $f$ restricted to $\overline{\phi}[h]$ is a non-zero map.
Proof.
Assume to the contrary that $f:\overline{\phi}[h]\rightarrow{\mathbb{F}}_{q^{m}}$ is the zero map. Then, $f$ is the zero map when restricted to $\overline{\phi}[T^{R}-a_{i}]$ for any $i\in\{1,\ldots,\ell\}$. Observe that for any $x\in\overline{\phi}[T^{R}-a_{i}]$, we have
(3.1) $f|_{\overline{\phi}[T^{R}-a_{i}]}(x)=\sum_{k=0}^{s}\left(g_{k}(a_{i}^{k}x)\right)=\sum_{k=0}^{s}a^{k}_{i}g_{k}(x).$
Note that the $q$-degree of the $q$-linearized polynomial is at most $r-1$, thus, if $\overline{\phi}[T^{R}-a_{i}]\subseteq\operatorname{Ker}\left(f|_{\overline{\phi}[T^{R}-a_{i}]}(x)\right)$, then, we get the following contradiction
$R=\operatorname{dim}_{{\mathbb{F}}_{q}}\overline{\phi}[T^{R}-a_{i}]\leq\operatorname{dim}_{{\mathbb{F}}_{q}}\operatorname{Ker}\left(f|_{\overline{\phi}[T^{R}-a_{i}]}\right)\leq r-1$
because $R=r+\delta-1$. Thus, $f|_{\overline{\phi}[T^{R}-a_{i}]}(x)$ must be the zero polynomial for any $k$. Consider the ×$\ell\times(s+1)$ Vandermonde matrix
$V(a_{1},\ldots,a_{\ell})=\begin{bmatrix}1&\cdots&a_{1}^{s}\\ \vdots&\ddots&\vdots\\ 1&\cdots&a_{\ell}^{s}\\ \end{bmatrix}$
Let $g_{k}(x)=a^{k}_{i}\sum_{j=0}^{r-1}u^{(k)}_{j}x^{[j]}$ in which $u^{(k)}_{j}\in{\mathbb{F}}_{q^{m}}$ for $10\leq j\leq r-$ 1. Rewrite Equation (3.1) to have
$\sum_{k=0}^{s}\sum_{j=0}^{r-1}a_{i}^{k}u^{(k)}_{j}x^{[j]}=\sum_{j=0}^{r-1}\left(\sum_{k=0}^{s}a_{i}^{k}u^{(k)}_{j}\right)x^{[j]}.$
Observe that for some $10\leq j\leq r-$ 1, the nonzero vector $(u^{(0)}_{j},\ldots,u^{(s)}_{j})\in\left({\mathbb{F}}_{q^{m}}\right)^{s+1}$ is in the kernel of $V(a_{1},\ldots,a_{\ell})$. However, since all the $a_{i}^{\prime}s$ are distinct, then, the matrix has full rank equal to $s+1$ and $\ell\geq s+1$. This forces that the map $x\mapsto V(a_{1},\dots,a_{\ell})$ is injective, a contradiction. ∎
Theorem 3.4.
The code defined by $\operatorname{enc}(\mathcal{M})$ is an optimal
$\left(m\times\ell R,m(s+1)r,\ell R-Rs-r+1,r,\delta\right)$
rank-metric code.
Proof.
Let $W=\overline{\phi}\left[\prod_{i=1}^{\ell}T^{R}-a_{i}\right]$. To prove the length of the code is $\ell R$, first, we show that $W_{i}$’s are orthogonal. Observe that $\overline{\phi}[T^{R}-a_{i}]$ is an ${\mathbb{F}}_{q}[T]-$module. Given $\operatorname{gcd}(T^{R}-a_{1},\ldots,T^{R}-a_{\ell})=1$, then, by Proposition 2.9, we have
$W=\overline{\phi}\left[\prod_{i=1}^{\ell}T^{R}-a_{i}\right]=\prod_{i=1}^{\ell}\overline{\phi}[T^{R}-a_{i}]=\prod_{i=1}^{\ell}W_{i}.$
The previous equation holds under the assumption that $W_{i}$’s are ${\mathbb{F}}_{q}[T]$-modules. Now, if we consider $W_{i}$’s as ${\mathbb{F}}_{q}$-vector spaces and since $W_{j}\cap\bigoplus_{\scriptstyle\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell}W_{i}=\{0\}$, then, $W=\bigoplus_{i=1}^{\ell}W_{i}$. In particular, $\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}\subseteq{\mathbb{F}}_{q^{m}}$ is an ${\mathbb{F}}_{q}$-basis of $W$ and thus $\dim_{{\mathbb{F}}_{q}}(W)=|\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}|=\ell R$, this shows that the length of $\operatorname{enc}(\mathcal{M})$ (as a vector rank-metric code) is $\ell R$. By fixing an ${\mathbb{F}}_{q}$-basis for ${\mathbb{F}}_{q^{m}}$, each element in $\operatorname{enc}(\mathcal{M})$ can be seen as a matrix in ${\mathbb{F}}_{q}^{m\times\ell R}$.
We now show that the ${\mathbb{F}}_{q^{m}}$-dimension of the code is $(s+1)r$. Set $f(\tau)\in\mathcal{M}$. From the construction, $f(\tau)=\sum^{s}_{k=0}g_{k}(\tau){\overline{\phi}}_{T^{R}}^{k}$ for some $g_{k}\in{\mathbb{F}}_{q^{m}}\{\tau\}_{\leq r-1}$. Now, for every $k\in\{0,\ldots,s\}$, we have $g_{k}(\tau)=\sum_{j=0}^{r-1}a^{(k)}_{j}\tau^{j}$ where $a^{(k)}_{j}\in{\mathbb{F}}_{q^{m}}$ for every $j\in\{0,\ldots,r-1\}$. This means that every vector
[…]
can be associated with $g_{0}(\tau),\ldots,g_{s}(\tau)\in{\mathbb{F}}_{q^{m}}\{\tau\}_{\leq r-1}$ and thus each $f(\tau)\in\mathcal{M}$. This shows that the ${\mathbb{F}}_{q^{m}}$-dimension of the code is $(s+1)r$, as long as the map $\operatorname{enc}$ is injective. To see the injectivity of $\operatorname{enc}$ simply observe that a polynomial of linearized degree $sR+r-1<(s+1)R\leq\ell R$ cannot be the zero map on a space of ${\mathbb{F}}_{q^{m}}$-dimension $\ell R$. This shows that the ${\mathbb{F}}_{q}$-dimension is $m(s+1)r$.
We now need to establish $(r,\delta)$-rank-locality according to Definition 2.28. Considering the code as a matrix rank-metric code (i.e. codewords are matrices), suppose $t\in\{1,\ldots,\ell R\}$ is a column index. Then, the $t$-th column in every codeword is an entry of the form $f(\beta^{(i_{t})}_{j_{t}})$ for some $i_{t}\in\{1,\ldots,\ell\}$ and $j_{t}\in\{1,\ldots,R\}$ where $\beta^{(i_{t})}_{j_{t}}$ is an element of the ${\mathbb{F}}_{q}$-basis of $W_{i_{t}}$. If $\{\beta^{(i_{t})}_{1},\ldots,\beta^{(i_{t})}_{R}\}$ is the ${\mathbb{F}}_{q}$-basis of $W_{i_{t}}$, then we define the set of column indexes $\Gamma(t)\subset\{1,\ldots,\ell R\}$ to be the set of indexes of the columns $f(\beta^{(i_{t})}_{1}),\ldots,f(\beta^{(i_{t})}_{R})$. It is obvious that $t\in\Gamma(t)$ and $|\Gamma(t)|=\dim_{{\mathbb{F}}_{q}}(W_{i_{t}})=R=r+\delta-1$, satisfying the first two conditions of Definition 2.28.
To check the third, consider the local code $\operatorname{enc}(\mathcal{M})|_{\Gamma(t)}$ associated with the $t$-th column. For a fixed $1\leq i_{t}\leq\ell$ and basis $\{\beta_{j}^{(i_{t})}\}_{1\leq j\leq R}\subseteq{\mathbb{F}}_{q^{m}}$ of $W_{i_{t}}=\overline{\phi}[T^{R}-a_{i_{t}}]$, the local code associated with the $t$-th column is as follows.
$\operatorname{enc}(\mathcal{M})|_{\Gamma(t)}=\left\{\left(f(\beta^{(i_{t})}_{1}),\ldots,f(\beta^{(i_{t})}_{R})\right)\bigg{|}f\in\mathcal{M}\right\}$
However, by Equation (3.1), we can rewrite $f(x)$ to be
$f|_{\overline{\phi}[T^{R}-a_{i_{t}}]}(x)=\sum_{k=0}^{s}a^{k}_{i_{t}}g_{k}(x)$
where $g_{k}(x)\in{\mathbb{F}}_{q^{m}}\langle x\rangle$ with $q$-degree at most $r-1$. If ${\mathbb{F}}_{q^{m}}\langle x\rangle_{\leq r-1}$ denotes the set of $q$-linearized polynomials with coefficients in ${\mathbb{F}}_{q^{m}}$ and $q$-degree at most $r-1$, then every element in ${\mathbb{F}}_{q^{m}}\langle x\rangle_{\leq r-1}$ can be written as $f|_{\overline{\phi}[T^{R}-a_{i_{t}}]}(x)$ for some $f\in\mathcal{M}$. In light of this, we observe that
$\operatorname{enc}(\mathcal{M})|_{\Gamma(t)}=\left\{\left(h(\beta^{(i_{t})}_{1}),\ldots,h(\beta^{(i_{t})}_{R})\right)\bigg{|}h\in{\mathbb{F}}_{q^{m}}\langle x\rangle_{\leq r-1}\right\}=:G.$
By Definition 2.27, $G$ is an $(R,r)$-Gabidulin code. Since Gabidulin codes are MRD and using Proposition 2.25, $d_{R}(\operatorname{enc}(\mathcal{M})|_{\Gamma(t)})=R-r+1=\delta$. This proves that $\operatorname{enc}(\mathcal{M})$ is an $\left(m\times\ell R,m(s+1)r\right)$ rank-metric code with $(r,\delta)$-rank-locality. In view of this and by Proposition 2.29, we conclude that
$d_{R}(\operatorname{enc}(\mathcal{M}))\leq\ell R-(s+1)r+1-(s+1-1)(R-r)=\ell R-Rs-r+1.$
Finally, observe that $\operatorname{enc}(\mathcal{M})$ is the set of evaluations of $q$-polynomials of the form $\sum^{s}_{k=0}g_{k}(x)\circ{\overline{\phi}}_{T^{R}}^{k}(x)$ with $q$-degree at most $Rs+r-1$ (maximum $q$-degree of $g_{k}(x)$ and $\overline{\phi}_{T^{R}}^{k}(x)$ are at most $r-1$ and $Rs$ respectively) over an ${\mathbb{F}}_{q}$-linearly independent set of size $\ell R\leq m$ which is $\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}$. In other words, $\operatorname{enc}(\mathcal{M})$ is a subcode of an $(\ell R,Rs+r)$ Gabidulin code. Hence
$d_{R}\left(\operatorname{enc}(\mathcal{M})\right)\geq\ell R-Rs-r+1.$
In conclusion, $\operatorname{enc}(\mathcal{M})$ is an optimal $\left(m\times\ell R,m(s+1)r,\ell R-Rs-r+1,r,\delta\right)$ rank-metric code.
4. Examples
4.1. Explicit Construction of a Rank Metric Code with Rank Locality
In this subsection we illustrate our construction with an example, for simplicity we directly work with $q$-linearized polynomials instead of polynomials in $\tau$.
Let us choose the following parameters: $q=5,r=2,\delta=2,\ell=3,s=2$. Thus, note that $R=r+\delta-1=3$. Take then $a_{1}=1,a_{2}=2,a_{3}=3$ and construct the polynomial $h\in{\mathbb{F}}_{5}[T]$ as:
$h=(T^{3}-1)(T^{3}-2)(T^{3}-3)$
Now choose $P\in{\mathbb{F}}_{5}[T]$ such as
$P=T^{10}+4T^{9}+4T^{7}+T^{6}+T^{4}+4T^{3}+4T+2$
A simple check shows that $P$ is irreducible and $P\equiv 1\pmod{h}$ (in particular it results $P=(T+4)h+1$). Let now ${\mathbb{F}}_{5}(\alpha)\cong{\mathbb{F}}_{5^{10}}$ be the splitting field of $P$ over ${\mathbb{F}}_{5}$ where $\alpha$ is a primitive element for the extension (with minimal polynomial $x^{10}+3x^{5}+3x^{4}+2x^{3}+4x^{2}+x+2$). Obviously $m=[{\mathbb{F}}_{5}(\alpha):{\mathbb{F}}_{5}]=10$ and notice that $R=3\nmid 10=m$, which implies that this code is not covered by the construction in [14].
Choose now $z\in{\mathbb{F}}_{5}(\alpha)$ to be a root of $P$, for example
$z=3\alpha^{8}+2\alpha^{7}+\alpha^{5}+4\alpha^{4}+3\alpha^{3}+2\alpha^{2}+3\alpha+2$
and construct the Carlitz Module
$\overline{\phi}_{T}(x)=zx+x^{5}.$
Therefore we can compute
$\overline{\phi}_{T^{3}}(x)=\overline{\phi}_{T}(\overline{\phi}_{T}(\overline{\phi}_{T}(x)))=x^{5^{3}}+(z^{25}+z^{5}+z)x^{5^{2}}+(z^{10}+z^{6}+z^{2})x^{5}+z^{3}x.$
Now we can construct $W_{1},W_{2},W_{3}$. In this example we have that each $W_{a_{i}}$ is a vector sub-space of ${\mathbb{F}}_{5}(\alpha)$, each one of dimension $R=3$ and ${\mathbb{F}}_{q}$-generated by the roots of $\overline{\phi}_{T^{3}}(x)-a_{i}x$, so that $\operatorname{Ker}(\phi_{T^{3}}(x)-a_{i}x)$, for $i\in\{1,2,3\}$. Thus, each of them is given by
$W_{1}=\langle\beta_{1}^{1},\beta_{2}^{1},\beta_{3}^{1}\rangle\quad W_{2}=\langle\beta_{1}^{2},\beta_{2}^{2},\beta_{3}^{2}\rangle\quad W_{3}=\langle\beta_{1}^{3},\beta_{2}^{3},\beta_{3}^{3}\rangle$
where we can choose
$\beta_{1}^{1}$ $=4\alpha^{8}+2\alpha^{7}+2\alpha^{5}+2\alpha^{4}+3\alpha^{2}+1$
$\beta_{2}^{1}$ $=\alpha^{9}+3\alpha^{8}+4\alpha^{7}+2\alpha^{6}+2\alpha^{5}+2\alpha^{4}+4\alpha^{2}+\alpha$
$\beta_{3}^{1}$ $=4\alpha^{9}+3\alpha^{8}+4\alpha^{7}+2\alpha^{5}+\alpha^{3}$
$\beta_{1}^{2}$ $=2\alpha^{9}+3\alpha^{8}+4\alpha^{7}+\alpha^{6}+3\alpha^{5}+\alpha^{4}+3\alpha^{3}+1$
$\beta_{2}^{2}$ $=4\alpha^{8}+\alpha^{7}+4\alpha^{6}+2\alpha^{5}+4\alpha^{4}+\alpha^{3}+\alpha$
$\beta_{3}^{2}$ $=4\alpha^{8}+3\alpha^{7}+\alpha^{6}+4\alpha^{3}+\alpha^{2}$
$\beta_{1}^{3}$ $=4\alpha^{7}+3\alpha^{6}+\alpha^{5}+2\alpha^{4}+2\alpha^{3}+1$
$\beta_{2}^{3}$ $=4\alpha^{9}+\alpha^{8}+4\alpha^{7}+4\alpha^{6}+2\alpha^{4}+\alpha$
$\beta_{3}^{3}$ $=\alpha^{9}+3\alpha^{7}+2\alpha^{6}+2\alpha^{5}+\alpha^{4}+2\alpha^{3}+\alpha^{2}$
The message space $\mathcal{M}$ for our code is
=$\mathcal{M}=\left\{\sum^{2}_{i=0}g_{i}(\overline{\phi}_{T^{3}}^{i}(x))\mid g_{i}\in{\mathbb{F}}_{5}(\alpha)[x],\deg_{q}(g_{i})\leq 1\right\}$
Now take for example $f\in\mathcal{M}$ such that $g_{0}=x+\alpha x^{5}$, $g_{1}=(\alpha+1)x^{5}$ and $g_{2}=(\alpha^{2}+\alpha)x$, namely:
$f=g_{0}(x)+g_{1}(\overline{\phi}_{T^{3}}(x))+g_{2}(\overline{\phi}_{T^{3}}(\overline{\phi}_{T^{3}}(x))).$
Then $\operatorname{enc}(f)=[f_{1},f_{2},f_{3},f_{4},f_{5},f_{6},f_{7},f_{8},f_{9}]$ where
$f_{1}$ $=f(\beta_{1}^{1})=2\alpha^{9}+\alpha^{8}+2\alpha^{7}+\alpha^{6}+3\alpha^{3}+3\alpha^{2}+4\alpha$
$f_{2}$ $=f(\beta_{2}^{1})=4\alpha^{9}+2\alpha^{8}+2\alpha^{6}+4\alpha^{5}+2\alpha^{4}+3\alpha^{3}+3\alpha^{2}+3\alpha+1$
$f_{3}$ $=f(\beta_{3}^{1})=3\alpha^{9}+3\alpha^{8}+\alpha^{7}+3\alpha^{6}+2\alpha^{5}+3\alpha^{4}+3\alpha^{2}+1$
$f_{4}$ $=f(\beta_{1}^{2})=3\alpha^{9}+\alpha^{8}+2\alpha^{6}+\alpha^{5}+\alpha^{4}+2\alpha^{2}+3\alpha+1$
$f_{5}$ $=f(\beta_{2}^{2})=3\alpha^{7}+3\alpha^{6}+3\alpha^{5}+2\alpha^{4}+4\alpha^{2}+4\alpha+2$
$f_{6}$ $=f(\beta_{3}^{2})=3\alpha^{9}+4\alpha^{8}+3\alpha^{7}+3\alpha^{6}+3\alpha^{5}+4\alpha^{4}+1$
$f_{7}$ $=f(\beta_{1}^{3})=4\alpha^{8}+2\alpha^{7}+2\alpha^{6}+3\alpha^{2}+\alpha+3$
$f_{8}$ $=f(\beta_{2}^{3})=2\alpha^{8}+3\alpha^{7}+2\alpha^{6}+2\alpha^{4}+\alpha^{3}+2\alpha^{2}+3$
$f_{9}$ $=f(\beta_{3}^{3})=2\alpha^{9}+2\alpha^{8}+2\alpha^{7}+2\alpha^{6}+4\alpha^{5}+\alpha^{4}+4\alpha^{3}+\alpha^{2}+\alpha+1$
Suppose now, for example, that $f_{5}$ is erased and we want to recover it. Since $W_{2}$ is a locality set, we can use $f_{4}$ and $f_{6}$ to construct a recovery polynomial $f|_{W_{2}}$ of $q$-degree at most $1$. Such polynomial is
$f|_{W_{2}}=(3\alpha+2)x^{5}+(4\alpha^{2}+4\alpha+1)x$
and it is easy to verify that $f|_{W_{2}}(\beta_{2}^{2})=f_{5}$.
4.2. Choosing the polynomial $P$
In our construction, a fundamental role is played by the polynomial $P$ that, a priori, might not exist. In Section 5 we show an infinite family of parameters for which $P$ exists. In addition, our polynomial exists also for many other choices of parameters that are not necessarily prescribed by Theorem 1.1: in the Table 1 we have reported many possible choices of $P=uh+1$ for some fixed values $q,R,\ell$ and points $a_{1},a_{2},a_{3}$ in ${\mathbb{F}}_{q}$.
5. An Infinite Family of Optimal Rank Metric Codes with Locality
5.1. Overview
In this section we will prove Theorem 1.1 as a corollary of our construction in Section 3 and Dirichlet Theorem for polynomial arithmetic progressions. We follow Rosen’s line of work [20] to extract the information needed to guarantee the existence of a polynomial of the wanted form. One of our tasks will be to precisely compute some error terms in [20, Chapter 4] (which is interesting on its own right, since effective versions of these results are difficult to find in the litarature).
5.2. Explicit Dirichlet Theorem
Let $\mathbb{C}$ be the field of complex numbers, $q$ be a prime power, $A={\mathbb{F}}_{q}[T]$ be the ring of polynomials over the finite field with $q$ elements, and $h$ be a fixed element of $A$ of positive degree. Unless otherwise specified, we write $P$ to denote a monic irreducible polynomial in $A$, so a sum whose index is $P$ is meant to run only over irreducible polynomials.
In what follows, we will need the result below. To prove it, we adjust the proof of [20, Theorem 4.8] to our special case in order to extract explicit constants.
Theorem 5.1.
[…]
|#$\left|\#S_{m}(a,h)-\frac{q^{m}}{m\Phi(h)}\right|\leq\left(3T(m)+2\deg(h)\right)\frac{q^{\frac{m}{2}}}{m}.$
Proof.
Set $H=\deg(h)$ and consider the $L$-series $L(s,\chi)$. We want to express it as a product in two different ways and compare the coefficients. By Proposition 2.21, we know that $L(s,\chi)$ is a polynomial in $q^{-s}$ of degree at most $H-1$. Therefore, setting $u=q^{-s}$ we can write
(5.1) $L^{*}(u,\chi)\coloneqq\sum_{k=0}^{H-1}a_{k}(\chi)u^{k}=\prod_{i=1}^{H-1}(1-\alpha_{i}(\chi)u)$
where $\alpha_{i}(\chi)$ is the inverse of a complex root of $L^{*}(u,\chi)$ for every $i\in\{1,\ldots,H-1\}$. The second expression for $L^{*}(u,\chi)$ comes from rewriting the Euler product for $L(s,\chi)$ in terms of $u$. Using Equation (2.5) and reordering the factors we have
$L(s,\chi)=\prod_{P\nmid h}\left(1-\frac{\chi(P)}{|P|^{s}}\right)^{-1}=\prod_{d=1}^{\infty}\prod_{\begin{subarray}{c}P\nmid h\\ \deg(P)=d\end{subarray}}(1-\chi(P)q^{-ds})^{-1}$
and setting again $u=q^{-s}$ we get
(5.2) $L^{*}(u,\chi)=\prod_{d=1}^{\infty}\prod_{\begin{subarray}{c}P\nmid h\\ \deg(P)=d\end{subarray}}(1-\chi(P)u^{d})^{-1}.$
The idea of the proof is to take the logarithmic derivative of both sides and compare the coefficients. We now state an identity for the logarithmic derivative which will be used frequently.
(5.3) $u\frac{d}{du}(\log(1-\alpha u)^{-1})=\sum_{m=1}^{\infty}\alpha^{m}u^{m}$
where $\alpha$ is a complex number. The proof is trivial using geometric series for $|u|<|\alpha|^{-1}$. Now, for each character $\chi$ modulo $h$ we define the number $c_{m}(\chi)$ such that
$u\frac{d}{du}\log(L^{*}(u,\chi))=\sum_{m=1}^{\infty}c_{m}(\chi)u^{m}.$
We have two cases: $\chi=\chi_{o}$ and $\chi\neq\chi_{o}$.
First, from Equation (2.3) and (2.4), writing $u=q^{-s}$ as always, we get
$L^{*}(u,\chi_{o})=\prod_{P|h}\left(1-u^{\deg(P)}\right)\frac{1}{1-qu}.$
Now, consider the inverse of the last equation. We take the logarithmic derivative and apply its additive property, we use Equation (5.3) and get
$u\frac{d}{du}\log(L^{*}(u,\chi_{o}))$ $=-u\frac{d}{du}\log(L^{*}(u,\chi_{o})^{-1})$
$=\sum_{m=1}^{\infty}q^{m}u^{m}-\sum_{m=1}^{\infty}\sum_{P|h}u^{m\deg(P)}$
$=\sum_{m=1}^{\infty}\left(q^{m}u^{m}-\sum_{P|h}u^{m\deg(P)}\right).$
Therefore, since in the worst case scenario for some $m$ the inner sum contributes with $-Hu^{m}$, comparing with the definition of $c_{m}(\chi)$, it is now obvious that
(5.4) $|c_{m}(\chi_{o})-q^{m}|\leq H.$
Instead, for $\chi\neq\chi_{o}$, take the logarithmic derivative of the inverse of $L^{*}(u,\chi)$ using Equation (5.1), multiply by $u$, and combine the resulting equation with Equation (5.3). We have
$c_{m}(\chi)=-\sum_{k=1}^{H-1}\alpha_{k}(\chi)^{m}.$
From the analogue of the Riemann hypothesis for function fields over a finite field, it follows that each of the $\alpha_{k}(\chi)$ has absolute value either 1 or $\sqrt{q}$ (see [2]). Therefore
(5.5) $|c_{m}(\chi)|\leq(H-1)q^{\frac{m}{2}}.$
Consider now the Euler product expansion of $L^{*}(u,\chi)$ given by Equation (5.2). Take the logarithmic derivative, multiply both sides of the resulting equation by $u$ and again, using Equation (5.3), we get
$c_{m}(\chi)=\sum_{\begin{subarray}{c}k,P\\ k\deg(P)=m\end{subarray}}\deg(P)\chi(P)^{k}$
that can be rearranged by separating the term corresponding to $k=1$ as outlined below.
$c_{m}(\chi)=m\sum_{\deg(P)=m}\chi(P)+\sum_{\begin{subarray}{c}d|m\\ d\leq m/2\end{subarray}}d\sum_{\deg(P)=d}\chi(P)^{m/d}.$
Since the value of a character is either zero or a root of unity, the absolute value of the inner sum on the right is less than or equal to the number of monic irreducible polynomials of degree $d$. Thus, using Theorem 2.15 we have
(5.6) $\left|c_{m}(\chi)-m\sum_{\deg(P)=m}\chi(P)\right|\leq\sum_{\begin{subarray}{c}d|m\\ d\leq m/2\end{subarray}}d\left(\frac{q^{d}}{d}+2\frac{q^{\frac{d}{2}}}{d}\right)\leq 3q^{\frac{m}{2}}T(m)$
where $T(m)$ is the number of divisors of $m$.
[…]
$\begin{split}&\left|\frac{1}{\Phi(h)}\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-m\#S_{m}(a,h)\right|=\\ &=\left|\frac{1}{\Phi(h)}\sum_{\chi}\bar{\chi}(a)\left(c_{m}(\chi)-m\sum_{\deg(P)=m}\chi(P)+m\sum_{\deg(P)=m}\chi(P)\right)-m\#S_{m}(a,h)\right|\\ &\leq U+V\end{split}$
where
$U=\left|\frac{1}{\Phi(h)}\sum_{\chi}\bar{\chi}(a)\left(c_{m}(\chi)-m\sum_{\deg(P)=m}\chi(P)\right)\right|$
and
$V=\left|\frac{m}{\Phi(h)}\sum_{\deg(P)=m}\sum_{\chi}\bar{\chi}(a)\chi(P)-m\#S_{m}(a,h)\right|.$
Finally, using Equation (5.6) for $U$, and the orthogonality relations from Proposition 2.18 for $V$, we can upper bound $U+V$ as follows:
$\begin{split}&U+V\leq\frac{3q^{\frac{m}{2}}}{\Phi(h)}T(m)\sum_{\chi}\left|\bar{\chi}(a)\right|+\left|m\sum_{\deg(P)=m}\delta(a,P)-m\#S_{m}(a,h)\right|\\ &=\frac{3q^{\frac{m}{2}}}{\Phi(h)}T(m)\Phi(h)=3q^{\frac{m}{2}}T(m)\end{split}$
namely
(5.7) $\left|\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-m\Phi(h)\#S_{m}(a,h)\right|\leq 3q^{\frac{m}{2}}T(m)\Phi(h).$
Next, using Equations (5.4) and (5.5), that $|\chi(a)|=1$ and $\chi_{o}(a)=1$ (since $\gcd(a,h)=1$), and the triangle inequality we get
(5.8) $\begin{split}\left|\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-q^{m}\right|&=\left|\bar{\chi_{o}}(a)c_{m}(\chi_{o})-q^{m}+\sum_{\chi\neq\chi_{o}}\bar{\chi}(a)c_{m}(\chi)\right|\\ &\leq H+\sum_{\chi\neq\chi_{o}}(H-1)q^{\frac{m}{2}}\leq 2Hq^{\frac{m}{2}}\Phi(h).\end{split}$
Finally, using the last two results (5.7) and (5.8) we have
(5.9) $\begin{split}\left|m\Phi(h)\#S_{m}(a,h)-q^{m}\right|=&\left|m\Phi(h)\#S_{m}(a,h)-\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)+\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-q^{m}\right|\\ \leq&\left|\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-m\Phi(h)\#S_{m}(a,h)\right|+\left|\sum_{\chi}\bar{\chi}(a)c_{m}(\chi)-q^{m}\right|\\ \leq&3q^{\frac{m}{2}}T(m)\Phi(h)+2Hq^{\frac{m}{2}}\Phi(h).\end{split}$
Dividing by $m\Phi(h)$ concludes the proof. ∎
5.3. Proof of the main theorem
We are now ready to easily prove the key part that will allow to construct the codes in Theorem 1.1.
Lemma 5.2.
Let $m\geq 4$ be an integer such that
$\ell R<\frac{m}{2}-\frac{\log(m)}{\log(q)}-\frac{\log\left(5\right)}{\log(q)}.$
Then there exists a monic irreducible polynomial $P\in{\mathbb{F}}_{q}[T]$ of degree $m$ such that $P\equiv 1\pmod{h}$.
Proof.
For $a=1$ and $h=\prod^{\ell}_{i=1}(T^{R}-a_{i})$ in Theorem 5.1 we know that
|#$\left|\#S_{m}(1,h)-\frac{q^{m}}{m\Phi(h)}\right|\leq 3\frac{q^{\frac{m}{2}}}{m}T(m)+2\deg(h)\frac{q^{\frac{m}{2}}}{m}.$
Since we can bound $T(m)\leq 2\sqrt{m}$ and $\deg(h)=\ell R\leq m$, we can write
#$\#S_{m}(1,h)\geq\frac{q^{m}}{m\Phi(h)}-6\frac{q^{\frac{m}{2}}}{\sqrt{m}}-2q^{\frac{m}{2}}.$
Bounding now $\Phi(h)$ simply with $q^{\ell R}$ and observing that we want to estimate $m$ for which #$\#S_{m}(1,h)$ is at least $1$, it is sufficient to check when
$0<\frac{q^{m}}{mq^{\ell R}}-6\frac{q^{\frac{m}{2}}}{\sqrt{m}}-2q^{\frac{m}{2}}.$
This gives
$\left(\frac{6}{\sqrt{m}}+2\right)<\frac{q^{\frac{m}{2}}}{mq^{\ell R}}.$
Taking the logarithm on both sides concludes the proof
$\ell R<\frac{m}{2}-\frac{\log(m)}{\log(q)}-\frac{\log\left(6/\sqrt{m}+2\right)}{\log(q)},$
since this condition is verified by hypothesis for $m\geq 4$. ∎
Finally, we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
With the framework we constructed, the proof of our result becomes immediate. First, since we have
$\ell R<\frac{m}{2}-\frac{\log(m)}{\log(q)}-\frac{\log\left(5\right)}{\log(q)},$
Lemma 5.2 guarantees the existence of an irreducible polynomial $P$ of degree $m$ that is congruent to $1$ modulo $h$. This allows for the construction of Section 3 to happen, as long as the conditions C1 to C4 are satisfied. Notice that C1,C2,C3 are directly satisfied by the hypothesis of Theorem 1.1, and C4 is forced by the stronger condition arising from Lemma 5.2 forcing the existence of $P$. This guarantees that the construction applies. Now derive the wanted parameters from Theorem 3.4. ∎
6. Acknowledgements
We would like to thank Federico Pellarin for interesting conversations on the topics of this paper. This work was supported by the National Science Foundation under Grant No 2338424.
Table 1.
[…]
5 3 3 $T^{9}+4T^{6}+T^{3}+4$ 10 $T+2$ $T^{10}+2T^{9}+4T^{7}+3T^{6}+T^{4}+2T^{3}+4T+4$
11 $T^{2}+T$ $T^{11}+T^{10}+4T^{8}+4T^{7}+T^{5}+T^{4}+4T^{2}+4T+1$
12 $T^{3}+2$ $T^{12}+T^{9}+4T^{6}+T^{3}+4$
13 $T^{4}+2T+4$ $T^{13}+T^{10}+4T^{9}+4T^{7}+T^{6}+T^{4}+4T^{3}+3T+2$
14 $T^{5}+4T$ $T^{14}+4T^{11}+4T^{10}+T^{8}+T^{7}+4T^{5}+4T^{4}+T+1$
$(a_{1},a_{2},a_{3})=(1,2,3)$ 15 $T^{6}+T+2$ $T^{15}+4T^{12}+T^{10}+3T^{9}+4T^{7}+2T^{6}+T^{4}+2T^{3}+4T+4$
16 $T^{7}+3$ $T^{16}+4T^{13}+T^{10}+3T^{9}+4T^{7}+2T^{6}+3T^{3}+3$
17 $T^{8}$ $T^{17}+4T^{14}+T^{11}+4T^{8}+1$
7 3 2 $T^{6}+4T^{3}+2$ 7 $T$ $T^{7}+4T^{4}+2T+1$
8 $T^{2}+1$ $T^{8}+T^{6}+4T^{5}+4T^{3}+2T^{2}+3$
9 $T^{3}+1$ $T^{9}+5T^{6}+6T^{3}+3$
$(a_{1},a_{2})=(1,2)$ 10 $T^{4}+T+5$ $T^{10}+5T^{7}+5T^{6}+6T^{4}+6T^{3}+2T+4$
11 $T^{5}+6$ $T^{11}+4T^{8}+6T^{6}+2T^{5}+3T^{3}+6$
8 3 2 $T^{6}+(\alpha+1)T^{3}+\alpha$ 7 $T+\alpha^{2}+\alpha$ $T^{7}+(\alpha^{2}+\alpha)T^{6}+(\alpha+1)T^{4}+T^{3}+\alpha T+\alpha^{2}+\alpha$
8 $T^{2}$ $T^{8}+(\alpha+1)T^{5}+\alpha T^{2}+1$
9 $T^{3}+\alpha T$ $T^{9}+\alpha T^{7}+(\alpha+1)T^{6}+(\alpha^{2}+\alpha)T^{4}+\alpha T^{3}+\alpha^{2}T+1$
$(a_{1},a_{2})=(1,\alpha)$ 10 $T^{4}$ $T^{10}+(\alpha+1)T^{7}+\alpha T^{4}+1$
11 $T^{5}$ $T^{11}+(\alpha+1)T^{8}+\alpha T^{5}+1$